# UAV-Enabled ISAC Systems with Fluid Antennas

### v1

@@ Abstract
Unmanned aerial vehicle (UAV)-enabled integrated sensing and communication (ISAC) is regarded as a key enabler for next-generation wireless systems. However, conventional fixed antenna arrays limit the ability of UAVs to fully exploit their inherent potential. To overcome this limitation, we propose a UAV-enabled ISAC framework equipped with fluid antenna (FA) arrays, where the mobility of antenna elements introduces additional spatial degrees of freedom to simultaneously enhance communication and sensing performance. A multi-objective optimization problem is formulated to maximize the communication rates of multiple users while minimizing the Cramér-Rao bound (CRB) for single-target angle estimation. Due to excessively frequent updates of FA positions may lead to response delays, a three-timescale optimization framework is developed to jointly design transmit beamforming, FA positions, and UAV trajectory based on their characteristics. To solve the non-convexity of the problem, an alternating optimization-based algorithm is developed to obtain a sub-optimal solution. Numerical results show that the proposed scheme significantly outperforms various benchmark schemes, validating the effectiveness of integrating the FA technology into the UAV-enabled ISAC systems.
@@ I Introduction
The emerging low-altitude economy (LAE) represents a transformative paradigm that leverages airspace within a few hundred meters above the ground to support a wide range of applications [1, 2]. The foundation of its development and future expansion lies in the deployment of robust communication infrastructures, particularly low-altitude wireless networks (LAWNs) [3]. Unlike conventional terrestrial networks, the LAWNs are specifically designed to meet the stringent requirements of aerial scenarios such as real-time monitoring, disaster response, and cargo delivery, offering ubiquitous connectivity with high reliability and low latency [4]. Unmanned aerial vehicles (UAVs), also referred to as drones or autonomous aerial vehicles (AAVs), are widely regarded as key enablers for the LAWNs. Owing to their operational efficiency and cost-effectiveness, the UAVs play a central role in a variety of the LAWN applications, thereby driving the rapid expansion of the LAE missions [5, 6]. However, the increasing scarcity of spectrum resources and the limited computation capabilities of the UAVs impose critical challenges on providing reliable and high-capacity communication services in the LAWNs, which in turn restrict sustainable growth of the LAE applications.
@@
Fortunately, fluid antennas (FAs), also known as movable antennas, have emerged as a promising solution to tackle these challenges [7, 8]. Unlike conventional fixed-position antennas (FPAs), the FAs dynamically adapt their physical positions in response to the time-varying dynamics of wireless channels [9, 10], which allows the system to opportunistically exploit the most favorable channel conditions in real time, thereby significantly improving signal reception and transmission efficiency in highly dynamic and congested spectral environments [11]. One of the most notable advantages of the FAs lies in their ability to mitigate interference and combat severe fading [12], both of which are particularly prevalent in the UAV communications due to high mobility and frequent line-of-sight (LoS) blockages. By adjusting the positions of the FAs, the UAV equipped with FAs can maintain robust connectivity and achieve superior throughput and latency performance, thus supporting mission-critical LAE applications [13].
@@
However, as the LAE missions over the LAWNs become increasingly complex, there is a growing demand for highly accurate and real-time environmental sensing. Conventional approaches, which rely on communication-dedicated systems or separate sensing devices, often face challenges in providing sufficient sensing capabilities without incurring additional energy consumption or spectrum overhead [14]. Integrated sensing and communication (ISAC) has emerged as a promising solution, offering a unified framework that simultaneously supports communication service and environmental sensing [15, 16, 17]. By leveraging shared spectral resources and hardware platforms, ISAC enables UAVs to perform tasks such as target detection, trajectory tracking, obstacle avoidance, and remote sensing while maintaining reliable communication links [18, 19, 20]. The dual-functional operation allows the UAVs to communicate and sense concurrently within the same hardware, significantly improving operational efficiency and safety, without further taxing limited resources such as spectrum, energy, or payload capacity. Such capabilities are particularly valuable in crowded or unpredictable low-altitude airspace, where the LAWNs must support highly dynamic LAE missions.
@@
Motivated by the growing demand for dual-functional ISAC applications in the LAWNs and the unique advantages of the FA technology, we propose in this paper a novel UAV-enabled ISAC system empowered by FA arrays. By exploiting the dual mobility of the UAV trajectory and the FA positions, the proposed system establishes a highly adaptive ISAC framework. Since excessively frequent updates of FA positions may cause response delays, a three-timescale optimization framework is developed to jointly design transmit beamforming, FA positions, and UAV trajectory in accordance with their characteristics. Compared to the existing FPA-based ISAC systems, the FA-aided UAV-ISAC system involves dual-scale spatial optimization, i.e., the small-scale antenna position adjustment within the FA arrays to exploit local spatial channel variations, and the large-scale UAV trajectory optimization for achieving favorable propagation conditions across the service region. The key contributions of this paper are summarized as follows:
@@
  • We study a UAV-enabled ISAC system, where a single UAV equipped with transmit and receive FA arrays simultaneously serves multiple communication users and senses a target via uplink echo signals. A multi-objective optimization problem is formulated to jointly maximize the multiuser communication rate and minimize the Cramér-Rao bound (CRB) for target angle estimation.
@@
  • To reduce the response delays caused by excessively frequent FA position updates, a three-timescale optimization framework is proposed to jointly design transmit beamforming, FA positions, and UAV trajectory according to their characteristics.
@@ II Related Works
II-A Fluid Antenna System
The fluid antenna system (FAS) has recently gained considerable attention due to its capability to dynamically switch antenna positions within a feasible region to select the optimal signal port. The reconfigurability enables remarkable diversity gain and resilience against fading, even in highly spatially correlated environments [12, 7]. Motivated by these advantages, several FAS-enabled frameworks have been proposed. For instance, a multiple access scheme exploiting the spatial reconfigurability of the FAS to achieve substantial performance improvements was proposed in [21], while [22] explored an uplink system where both the base station and user terminals were equipped with the FAs to maximize capacity. Moreover, the FAS has been integrated into the ISAC systems. In particular, [23] developed an FAS-assisted multiple-input and multiple-output (MIMO) ISAC framework to simultaneously enhance the communication quality and the radar sensing, and a two-dimensional FAS-aided ISAC system was developed in [24] for multiuser MIMO downlink networks that jointly optimized the FA positions and the precoding. In addition to these core studies, the integration of the FAS with other emerging technologies has also been explored. For example, the UAV relay-enabled systems incorporating the FAS with non-orthogonal multiple access (NOMA) were investigated in [25], and [26] examined the UAV-enabled multiuser communication where ground users were equipped with the FAS under a rate-splitting multiple access (RSMA) framework. Overall, these studies highlighted the broad applicability of the FAS in advancing the ISAC systems and further revealed its strong potential for the UAV-enabled communication paradigms.
II-B UAV-enabled ISAC System
@@ III System Model and Problem Formulation
As shown in Fig.  1, a dual-function radar-communication UAV is deployed to serve $M$ downlink users and detect a single-point target. Let $\mathcal{M}=\{1,\dots,M\}$ denote the set of users. Specifically, the UAV transmits ISAC signals to provide communication services to users while simultaneously leveraging the reflected echoes from the target for radar sensing. Assuming that each user is equipped with a single receive antenna, and the UAV is equipped with the FA arrays of sizes $N_{t}$ and $N_{r}$ at the transmitter and receiver sides, respectively.
Without loss of generality, a three-dimensional (3D) Cartesian coordinate system is considered. The UAV is assumed to fly at a fixed altitude $H$, starting from a given initial location $\mathbf{q}_{\rm{I}}$ and ending at a final location $\mathbf{q}_{\rm{F}}$. For analytical tractability, the UAV’s flight duration is discretized into $N$ equal time slots, each of duration $\tau=T/N$, where $T$ denotes the total mission period. Let $\mathcal{N}=\{1,\dots,N\}$ denote the set of time slots. The horizontal location of the UAV at time slot $n$ is denoted by $\mathbf{q}[n]=(x[n],y[n])$, and its full 3D location is therefore $(\mathbf{q}[n],H)$. The users are assumed to be static, with user $m$ located at $(\mathbf{q}_{m},0)$, where $\mathbf{q}_{m}=(x_{m},y_{m})$ denotes its horizontal location. Similarly, the target is located at $(\mathbf{q}_{\rm{T}},0)$, where $\mathbf{q}_{\rm{T}}=(x_{\rm{T}},y_{\rm{T}})$ denotes its horizontal location. Furthermore, the positions of transmit and receive FA arrays at time slot $n$ are defined as $\mathbf{x}[n]=\{x_{1}[n],x_{2}[n],\ldots,x_{N_{t}}[n]\}^{\mathsf{T}}$ and $\mathbf{y}[n]=\{y_{1}[n],y_{2}[n],\ldots,y_{N_{r}}[n]\}^{\mathsf{T}}$, respectively  [^1] Due to the equivalence of the antennas, we can assume $x_{1}[n]<x_{2}[n]<\dots<x_{N_{t}}[n],\forall n\in\mathcal{N}$ and $y_{1}[n]<y_{2}[n]<\ldots<y_{N_{r}}[n],\forall n\in\mathcal{N}$..
Figure 1: Illustration of the UAV-enabled ISAC system with transmit and receive FA arrays vertically mounted on the UAV.
@@ III-A Channel Model
We consider that the channel vector is determined not only by the signal propagation environment but also by the positions of the FAs. The far-field channel model is adopted, where the size of the transmit region is considerably smaller than the signal propagation distance [25, 26]. Accordingly, all transmit FAs experience the same angle of departure (AoD) and the same amplitude of the complex path coefficient, while only the phase varies with the positions of the transmit FAs.
Let $\theta_{m}[n]$ denote the vertical AoD from the UAV to the user $m$ at time slot $n$, which is given by
@@
Then, the signal propagation phase difference between the $k$-th transmit FA position ($k=1,2,\dots,N_{t}$) and the original position at time slot $n$ is given by $\frac{2\pi}{\lambda}x_{k}[n]\sin(\theta_{m}[n])$, where $\lambda$ denotes the carrier wavelength. Therefore, the transmit array response vector from the UAV to the user $m$ is given by
@@
Next, the sensing channel model between the UAV and the target is considered. Following [24], a single-path propagation model is adopted in the considered ISAC system, as multipath reflections often suffer from significant attenuation in complex environments. Therefore, the steering vectors of the transmit and receive FA arrays can be respectively given by
@@
where $\theta_{\rm{T}}[n]$ denotes the vertical AoD from the UAV to the target at time slot $n$.
@@ III-B User Communication Model
We first consider the downlink communication from the UAV to the user $m$, where the UAV transmits an ISAC signal using coordinated transmit beamforming. [^3] In this paper, it is assumed that the UAV is equipped with a global positioning system (GPS) and can acquire the real-time location information of communication users through positioning systems. Accordingly, the large-scale fading is computed based on a geometric path-loss model, while the small-scale fading is characterized using distributions parameterized by the LoS probability. By combining these two components, the statistical channel state information (CSI) is generated for system design and performance analysis. Let $\tilde{L}>N_{t}$ denote the length of the transmission frame at time slot $n$, where $\tilde{l}\in\{1,2,\dots,\tilde{L}\}$ represents the discrete time index within the frame. Specifically, the transmitted signal from the UAV at time slot $n$ is given by
@@
where $\mathbf{w}_{m}[n]\in\mathbb{C}^{N_{t}\times 1}$ denotes the transmit beamforming vector directed toward user $m$, and $s_{m}^{\mathrm{c}}[\tilde{l},n]\sim\mathcal{CN}(0,1)$ represents the corresponding information symbol, assumed to follow the circularly symmetric complex Gaussian (CSCG) distribution. $\mathbf{s}_{0}[\tilde{l},n]$ denotes the dedicated sensing signal, which follows a zero-mean CSCG distribution with covariance matrix $\mathbf{R}_{0}[n]=\mathbb{E}\{\mathbf{s}_{0}[\tilde{l},n]\mathbf{s}^{\mathsf{H}}_{0}[\tilde{l},n]\}\succeq\mathbf{0}$.  [^4] It is noted that in (7), we consider a multi-beam transmission strategy for dedicated radar signals $\mathbf{s}_{0}[\tilde{l},n]$ [29]. Such a design can be directly extended to the detection of multiple targets. Consequently, $\mathbf{R}_{0}[n]$ is of general rank, with $\text{rank}(\mathbf{R}_{0}[n])=N_{\mathrm{s}}$, where $0\leq N_{\mathrm{s}}\leq N_{t}$. This consideration corresponds to forming a set of $N_{\mathrm{s}}$ radar signal beams in time slot $n$, which can be obtained through the eigenvalue decomposition of $\mathbf{R}_{0}[n]$.
@@
$z_{m}^{\mathrm{c}}[\tilde{l},n]=\underbrace{\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{m}[n]s_{m}^{\mathrm{c}}[\tilde{l},n]}_{\rm{Desired\ signal}}+\underbrace{\sum\limits_{i=1,i\neq m}^{\mathcal{M}}\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{i}[n]s_{i}^{\mathrm{c}}[\tilde{l},n]}_{\rm{Inter-user\ interference}}$
@@
$\gamma_{m}[n]=\frac{|\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{m}[n]|^{2}}{\sum\limits_{i=1,i\neq m}^{\mathcal{M}}|\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{i}[n]|^{2}+\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\mathbf{h}_{m}[n]+\sigma^{2}_{m}}.$ (9)

### v2

# UAV-Enabled ISAC with Fluid Antennas for Low-Altitude Wireless Networks

@@ Abstract
Unmanned aerial vehicle (UAV)-enabled integrated sensing and communication (ISAC) is regarded as a key enabler for next-generation wireless systems. However, conventional fixed-position antennas limit the ability of UAVs to fully exploit their inherent potential. To overcome this limitation, we propose a UAV-enabled ISAC framework equipped with fluid antennas (FAs), where the mobility of antenna elements introduces additional spatial degrees of freedom to simultaneously enhance communication and sensing performance. A multi-objective optimization problem is formulated to maximize the communication rates of multiple users while minimizing the Cramér-Rao bound (CRB) for the angle estimation of a single target. Due to excessively frequent updates of FA positions may lead to response delay, a three-timescale optimization framework is developed to jointly optimize transmit beamforming, FA positions, and UAV trajectory based on their characteristics. To solve the non-convexity of the problem, an alternating optimization-based algorithm is developed to obtain a sub-optimal solution. Numerical results show that the proposed scheme significantly outperforms various benchmark schemes, validating the effectiveness of integrating the FA technology into the UAV-enabled ISAC systems.
@@ I Introduction
The emerging low-altitude economy (LAE) represents a transformative paradigm that leverages airspace within a few hundred meters above the ground to support a wide range of applications [1, 2]. The foundation of its development and future expansion lies in the deployment of robust communication infrastructures, particularly low-altitude wireless networks (LAWNs) [3]. Unlike conventional terrestrial networks, the LAWNs are specifically designed to meet the stringent requirements of aerial scenarios such as real-time monitoring, disaster response, and cargo delivery, offering ubiquitous connectivity with high reliability and low latency [4]. Unmanned aerial vehicles (UAVs), also referred to as drones or autonomous aerial vehicles (AAVs), are widely regarded as a key enabler for the LAWNs. Owing to their operational efficiency and cost-effectiveness, the UAVs play a central role in a variety of the LAE applications, thereby driving the rapid expansion of the LAE missions [5, 6]. However, the increasing scarcity of spectrum resources and the limited computation capabilities of the UAVs impose critical challenges on providing reliable and high-capacity communication services in the LAWNs, which in turn restrict sustainable growth of the LAE applications.
@@
Fortunately, fluid antennas (FAs), also known as movable antennas, have emerged as a promising solution to tackle these challenges [7, 8]. Unlike conventional fixed-position antennas (FPAs), the FAs dynamically adapt their physical positions in response to the time-varying dynamics of wireless channels [9, 10], which allows the system to opportunistically exploit the most favorable channel conditions in real time, thereby significantly improving signal reception and transmission efficiency in highly dynamic and congested environments [11]. One of the most notable advantages of the FAs lies in their ability to mitigate interference and combat severe fading [12], both of which are particularly prevalent in the UAV communications due to high mobility and frequent line-of-sight (LoS) blockages. By adjusting the FA positions, the UAV equipped with FAs can maintain robust connectivity and achieve superior throughput and latency performance, thus supporting mission-critical LAE applications [13].
@@
However, as the LAE missions over the LAWNs become increasingly complex, there is a growing demand for highly accurate and real-time environmental sensing. Conventional approaches, which rely on communication-dedicated systems or separate sensing devices, often face challenges in providing sufficient sensing capabilities without incurring additional energy consumption or spectrum overhead [14]. Integrated sensing and communication (ISAC) has emerged as a promising solution, offering a unified framework that simultaneously supports communication services and environmental sensing [15, 16, 17]. By leveraging shared spectral resources and hardware platforms, ISAC enables UAVs to perform tasks such as target detection, trajectory tracking, obstacle avoidance, and remote sensing while maintaining reliable communication links [18, 19, 20]. The dual-functional operation allows the UAVs to communicate and sense concurrently within the same hardware, significantly improving operational efficiency and safety, without further taxing limited resources such as spectrum, energy, or payload capacity. Such capabilities are particularly valuable in crowded or unpredictable low-altitude airspace, where the LAWNs must support highly dynamic LAE missions.
@@
Motivated by the growing demand for dual-functional ISAC applications in the LAWNs and the unique advantages of the FA technology, we propose in this paper a novel UAV-enabled ISAC system with FAs. By exploiting the dual mobility of the UAV trajectory and the FA positions, the proposed system establishes a highly adaptive ISAC framework. Since excessively frequent updates of FA positions may cause response delay, a three-timescale optimization framework is developed to jointly design transmit beamforming, FA positions, and UAV trajectory in accordance with their characteristics. Compared to the existing FPA-based ISAC systems, the FA-aided UAV-ISAC system involves dual-scale spatial optimization, i.e., the small-scale antenna position adjustment within the FA arrays to exploit local spatial channel variations, and the large-scale UAV trajectory optimization for achieving favorable propagation conditions across the service region. The key contributions of this paper are summarized as follows:
@@
  • We study a UAV-enabled ISAC system, where a single UAV equipped with transmit and receive FAs simultaneously serves multiple communication users and senses a single target via uplink echo signals. A multi-objective optimization problem is formulated to jointly maximize the multiuser communication rate and minimize the Cramér-Rao bound (CRB) for target angle estimation.
@@
  • To reduce the response delay caused by excessively frequent FA position updates, a three-timescale optimization framework is proposed to jointly optimize transmit beamforming, FA positions, and UAV trajectory according to their characteristics.
@@ II Related Works
The rapid development of UAV-enabled ISAC systems has created new opportunities for flexible wireless system design. Meanwhile, FA systems (FASs) have attracted growing attention due to their ability to dynamically adjust antenna positions within a confined region, thereby providing additional spatial degrees of freedom (DoFs). Although these two research directions have been studied individually, their integration remains largely unexplored. This section therefore reviews the most relevant studies on FASs and UAV-enabled ISAC systems.
II-A Fluid Antenna Systems
The FAS has recently gained considerable attention due to its capability to dynamically adjust antenna positions within a feasible region to select the optimal signal port. The reconfigurability enables remarkable diversity gain and resilience against fading, even in highly spatially correlated environments [12, 7]. Motivated by these advantages, several FAS-enabled frameworks have been proposed. For instance, a multiple access scheme exploiting the spatial reconfigurability of the FAS to achieve substantial performance improvements was proposed in [21], while [22] explored an uplink system where both the base station and user terminals were equipped with the FAs to maximize channel capacity. Moreover, the FAS has been integrated into the ISAC systems. In particular, [23] developed an FAS-assisted multiple-input and multiple-output (MIMO) ISAC framework to simultaneously enhance the communication quality and the radar sensing, and a two-dimensional FAS-aided ISAC system was developed in [24] for multiuser MIMO downlink networks that jointly optimized the FA positions and the precoding. In addition to these core studies, the integration of the FAS with other emerging technologies has also been explored. For example, the UAV relay-enabled systems incorporating the FAS with non-orthogonal multiple access (NOMA) were investigated in [25], and [26] examined the UAV-enabled multiuser communication where ground users were equipped with the FAS under a rate-splitting multiple access (RSMA) framework. Overall, these studies highlighted the broad applicability of the FAS in advancing the ISAC systems and further revealed its strong potential for enhancing the UAV-enabled communication paradigms.
II-B UAV-enabled ISAC Systems
@@ III System Model and Problem Formulation
As shown in Fig.  1, a dual-function radar-communication UAV is deployed to serve $M$ downlink users and detect a single-point target. Let $\mathcal{M}=\{1,\dots,M\}$ denote the set of users. Specifically, the UAV transmits ISAC signals to provide communication services to users while simultaneously leveraging the reflected echoes from the target for radar sensing. Assuming that each user is equipped with a single antenna, and the UAV is equipped with two FA arrays of sizes $N_{t}$ and $N_{r}$ at the transmitter and receiver sides, respectively.
Without loss of generality, a three-dimensional (3D) Cartesian coordinate system is considered. The UAV is assumed to fly at a fixed altitude $H$, starting from a given initial location $\mathbf{q}_{\rm{I}}$ and ending at a final location $\mathbf{q}_{\rm{F}}$. For analytical tractability, the UAV’s flight duration is discretized into $N$ equal time slots, each of duration $\tau=T/N$, where $T$ denotes the total mission period. Let $\mathcal{N}=\{1,\dots,N\}$ denote the set of time slots. The horizontal location of the UAV in time slot $n$ is denoted by $\mathbf{q}[n]=(x[n],y[n])$, and its full 3D location is therefore $(\mathbf{q}[n],H)$. The users are assumed to be static, with user $m$ located at $(\mathbf{q}_{m},0)$, where $\mathbf{q}_{m}=(x_{m},y_{m})$ denotes its horizontal location. Similarly, the target is located at $(\mathbf{q}_{\rm{T}},0)$, where $\mathbf{q}_{\rm{T}}=(x_{\rm{T}},y_{\rm{T}})$ denotes its horizontal location. Furthermore, the transmit and receive FA positions in time slot $n$ are defined as $\mathbf{x}[n]=\{x_{1}[n],x_{2}[n],\ldots,x_{N_{t}}[n]\}^{\mathsf{T}}$ and $\mathbf{y}[n]=\{y_{1}[n],y_{2}[n],\ldots,y_{N_{r}}[n]\}^{\mathsf{T}}$, respectively  [^1] Due to the equivalence of the antennas, we can assume $x_{1}[n]<x_{2}[n]<\dots<x_{N_{t}}[n],\forall n\in\mathcal{N}$ and $y_{1}[n]<y_{2}[n]<\ldots<y_{N_{r}}[n],\forall n\in\mathcal{N}$..
Figure 1: Illustration of the UAV-enabled ISAC system with transmit and receive FAs vertically mounted on the UAV.
@@ III-A Channel Model
We consider that the channel vector is determined not only by the signal propagation environment but also by the FA positions. The far-field channel model is adopted, where the size of the transmit region is considerably smaller than the signal propagation distance [25, 26]. Accordingly, all transmit FAs experience the same angle of departure (AoD) and the same amplitude of the complex path coefficient, while only the phase varies with their positions.
Let $\theta_{m}[n]$ denote the vertical AoD from the UAV to the user $m$ in time slot $n$, which is given by
@@
Then, the signal propagation phase difference between the $k$-th transmit FA position ($k=1,2,\dots,N_{t}$) and the original position in time slot $n$ is given by $\frac{2\pi}{\lambda}x_{k}[n]\sin(\theta_{m}[n])$, where $\lambda$ denotes the carrier wavelength. Therefore, the transmit array response vector from the UAV to the user $m$ is given by
@@
Next, the sensing channel model between the UAV and the target is considered. Following [24], a single-path propagation model is adopted in the considered ISAC system, as multipath reflections often suffer from significant attenuation in complex environments. Therefore, the steering vectors of the transmit and receive FAs can be respectively given by
@@
where $\theta_{\rm{T}}[n]$ denotes the vertical AoD from the UAV to the target in time slot $n$.
@@ III-B User Communication Model
We first consider the downlink communication from the UAV to the user $m$, where the UAV transmits an ISAC signal using coordinated transmit beamforming  [^3] In this paper, it is assumed that the UAV is equipped with a global positioning system (GPS) and can acquire the real-time location information of users through positioning systems. Accordingly, the large-scale fading is computed based on a geometric path-loss model, while the small-scale fading is characterized using distributions parameterized by the LoS probability. By combining these two components, the statistical channel state information (CSI) is generated for system design and performance analysis.. Let $\tilde{L}>N_{t}$ denote the length of the transmission frame in time slot $n$, where $\tilde{l}\in\{1,2,\dots,\tilde{L}\}$ represents the discrete time index within the frame. Specifically, the transmitted signal from the UAV in time slot $n$ is given by
@@
where $\mathbf{w}_{m}[n]\in\mathbb{C}^{N_{t}\times 1}$ denotes the transmit beamforming vector directed toward the user $m$, and $s_{m}^{\mathrm{c}}[\tilde{l},n]\sim\mathcal{CN}(0,1)$ represents the corresponding information symbol, assumed to follow the circularly symmetric complex Gaussian (CSCG) distribution. $\mathbf{s}_{0}[\tilde{l},n]$ denotes the dedicated sensing signal, which follows a zero-mean CSCG distribution with covariance matrix $\mathbf{R}_{0}[n]=\mathbb{E}\{\mathbf{s}_{0}[\tilde{l},n]\mathbf{s}^{\mathsf{H}}_{0}[\tilde{l},n]\}\succeq\mathbf{0}$.  [^4] It is noted that in (7), we consider a multi-beam transmission strategy for dedicated radar signals $\mathbf{s}_{0}[\tilde{l},n]$ [29]. Such a design can be directly extended to the detection of multiple targets. Consequently, $\mathbf{R}_{0}[n]$ is of general rank, with $\text{rank}(\mathbf{R}_{0}[n])=N_{\mathrm{s}}$, where $0\leq N_{\mathrm{s}}\leq N_{t}$. This consideration corresponds to forming a set of $N_{\mathrm{s}}$ radar signal beams in time slot $n$, which can be obtained through the eigenvalue decomposition of $\mathbf{R}_{0}[n]$.
@@
$z_{m}^{\mathrm{c}}[\tilde{l},n]=\underbrace{\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{m}[n]s_{m}^{\mathrm{c}}[\tilde{l},n]}_{\rm{Desired\ signal}}+\underbrace{\sum\limits_{i\neq m}^{\mathcal{M}}\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{i}[n]s_{i}^{\mathrm{c}}[\tilde{l},n]}_{\rm{Inter-user\ interference}}$
@@
$\gamma_{m}[n]=\frac{|\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{m}[n]|^{2}}{\sum\limits_{i\neq m}^{\mathcal{M}}|\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{w}_{i}[n]|^{2}+\mathbf{h}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\mathbf{h}_{m}[n]+\sigma^{2}_{m}}.$ (9)